the displacement of a body is given by 2s=gt^2,where g is constant,the vel of the body at time t is
a)gt b)gt/2 C)gt^2/2 D)gt^2/2
1
Expert's answer
2012-09-07T09:04:14-0400
s=gt^2/2
v= s’=( gt^2/2) = gt